### v1

# Faster Gradient Methods for Highly-smooth Stochastic Bilevel Optimization

Abstract
This paper studies the complexity of finding an $\epsilon$-stationary point for stochastic bilevel optimization when the upper-level problem is nonconvex and the lower-level problem is strongly convex. Recent work proposed the first-order method, F2SA, achieving the $\tilde{\mathcal{O}}(\epsilon^{-6})$ upper complexity bound for first-order smooth problems. This is slower than the optimal $\Omega(\epsilon^{-4})$ complexity lower bound in its single-level counterpart. In this work, we show that faster rates are achievable for higher-order smooth problems. We first reformulate F2SA as approximating the hyper-gradient with a forward difference. Based on this observation, we propose a class of methods F2SA-$p$ that uses $p$th-order finite difference for hyper-gradient approximation and improves the upper bound to $\tilde{\mathcal{O}}(p\epsilon^{4-p/2})$ for $p$th-order smooth problems. Finally, we demonstrate that the $\Omega(\epsilon^{-4})$ lower bound also holds for stochastic bilevel problems when the high-order smoothness holds for the lower-level variable, indicating that the upper bound of F2SA-$p$ is nearly optimal in the highly smooth region $p=\Omega(\log\epsilon^{-1}/\log\log\epsilon^{-1})$.
{NoHyper}
1 Introduction
Many machine learning problems, such as meta-learning [39], hyper-parameter tuning [5, 16, 36], and adversarial training [18] can be abstracted as solving the following bilevel optimization problem:
$\min_{{\bm{x}}\in{\mathbb{R}}^{d_{x}}}\varphi({\bm{x}})=f({\bm{x}},{\bm{y}}^{*}({\bm{x}})),\quad{\bm{y}}^{*}({\bm{x}})=\arg\min_{{\bm{y}}\in{\mathbb{R}}^{d_{y}}}g({\bm{x}},{\bm{y}}),$ (1)
We call $f$ and $g$ the upper-level and lower-level functions, respectively, and call $\varphi$ the hyper-objective. In this paper, we consider the most common nonconvex-strongly-convex setting where $f:{\mathbb{R}}^{d_{x}}\rightarrow{\mathbb{R}}$ is smooth and possibly nonconvex, and $g:{\mathbb{R}}^{d_{y}}\rightarrow{\mathbb{R}}$ is smooth jointly in $({\bm{x}},{\bm{y}})$ and strongly convex in $\bm{y}$}. Under the lower-level strong convexity assumption, the implicit function theorem indicates the following closed form of the hyper-gradient [17]:
$\nabla\varphi({\bm{x}})=\nabla_{x}f({\bm{x}},{\bm{y}}^{*}({\bm{x}}))-\nabla_{xy}^{2}g({\bm{x}},{\bm{y}}^{*}({\bm{x}}))[\nabla_{yy}^{2}g({\bm{x}},{\bm{y}}^{*}({\bm{x}}))]^{-1}\nabla_{y}f({\bm{x}},{\bm{y}}^{*}({\bm{x}})).$ (2)
Following the works in nonconvex optimization [6, 7, 3], we consider the task of finding an $\epsilon$-stationary point of $\varphi$, i.e., a point ${\bm{x}}\in{\mathbb{R}}^{d_{x}}$ such that $\|\nabla\varphi({\bm{x}})\|\leq~\epsilon$. Motivated by many real machine learning tasks, we study the stochastic setting, where the algorithms only have access to the stochastic derivative estimators of both $f$ and $g$.
The first efficient algorithm BSA Ghadimi and Wang [17] for solving the stochastic bilevel problem leverages both stochastic gradient and Hessian-vector-product (HVP) oracles to find an $\epsilon$-stationary point of $\varphi({\bm{x}})$. Subsequently, Ji et al. [21] proposed stocBiO by incorporating multiple enhanced designs to improve the complexity. Both BSA and stocBiO require the stochastic Hessian assumption (5) on the lower-level function, which means $g$ has an unbiased stochastic Hessian estimator with bounded variance. For finite-sum problems, such an assumption is stronger than standard SGD assumptions and equivalent to the mean-squared-smoothness assumption (6) on the lower-level gradient estimator $G$ [2, Observation 1 and 2].
To avoid estimating HVP oracles, Kwon et al. [28] proposed the first fully first-order method F2SA that works under standard SGD assumptions on both $f$ and $g$ (Assumption 2.1). The main idea is to solve the following penalty problem [33, 34, 41]:
$\min_{{\bm{x}}\in{\mathbb{R}}^{d_{x}},{\bm{y}}\in{\mathbb{R}}^{d_{y}}}f({\bm{x}},{\bm{y}})+\lambda\left(g({\bm{x}},{\bm{y}})-\min_{{\bm{z}}\in{\mathbb{R}}^{d_{y}}}g({\bm{x}},{\bm{z}})\right),$ (3)
where $\lambda$ is taken to be sufficiently large such that $\lambda=\Omega(\epsilon^{-1})$. Thanks to Danskin’s theorem, the gradient of the penalty function in Eq. (3) only involves gradient information. Therefore, F2SA does not require the stochastic Hessian assumptions (5). More importantly, by directly leveraging gradient oracles instead of more expensive HVP oracles, the F2SA is more efficient in practice [40, 42, 22] and it is also the only method that can be scaled to 3232B sized large language model (LLM) training [38].
Kwon et al. [28] proved that the F2SA method finds an $\epsilon$-stationary point of $\varphi({\bm{x}})$ with $\tilde{\mathcal{O}}(\epsilon^{-3})$ first-order oracle calls in the deterministic case and $\tilde{\mathcal{O}}(\epsilon^{-7})$ stochastic first-order oracle (SFO) calls in the stochastic case. Recently, Chen et al. [10] showed the two-time-scale stepsize strategy improves the upper complexity bound of F2SA method to $\tilde{\mathcal{O}}(\epsilon^{-2})$ in the deterministic case, which is optimal up to logarithmic factors. However, the direct extension of their method in the stochastic case leads to the $\tilde{\mathcal{O}}(\epsilon^{-6})$ SFO complexity [10, 29] , which still has a significant gap between the $\Omega(\epsilon^{-4})$ lower bound for SGD [3]. It remains open whether optimal rates for stochastic bilevel problems can be achieved for fully first-order methods.
In this work, we revisit F2SA and interpret it as using forward difference to approximate the hyper-gradient. Our novel interpretation in turn leads to straightforward algorithm extensions for the F2SA method. Observing that the forward difference used by F2SA only has a first-order error guarantee, a natural idea to improve the error guarantee is to use higher-order finite difference methods. For instance, we know that the central difference has an improved second-order error guarantee. Based on this fact, we can derive the F2SA-2 method that solves the following symmetric penalty problem:
$\min_{{\bm{x}}\in{\mathbb{R}}^{d_{x}},{\bm{y}}\in{\mathbb{R}}^{d_{y}}}\frac{1}{2}\left(f({\bm{x}},{\bm{y}})+\lambda g({\bm{x}},{\bm{y}})-\min_{{\bm{z}}\in{\mathbb{R}}^{d_{y}}}\left(-f({\bm{x}},{\bm{z}})+\lambda g({\bm{x}},{\bm{z}})\right)\right).$ (4)
We can show that F2SA-2 returns an $\epsilon$-estimation to $\nabla\varphi({\bm{x}})$ under the setting $\lambda=\Omega(\epsilon^{-1/2})$ instead of $\Omega(\epsilon^{-1})$ in F2SA, which further improves the SFO complexity of F2SA from $\tilde{\mathcal{O}}(\epsilon^{-6})$ to $\tilde{\mathcal{O}}(\epsilon^{-5})$ for second-order smooth problems. Our idea is generalizable for any $p$th-order problems. We recall that in numerical analysis there exists the $p$th-order central difference that uses $p$ points to construct an estimator to the derivative of a unitary function with $p$th-order error guarantee (Lemma 3.1). Motivated by this fact, we propose the F2SA-$p$ algorithm and show that it allows $\lambda=\Omega(\epsilon^{-1/p})$ for $p$th-order smooth problems, which further leads to the improved $\tilde{\mathcal{O}}(p\epsilon^{-4-2/p})$ SFO complexity for finding an $\epsilon$-stationary point stated by our Theorem 3.1.
To examine the tightness of our upper bounds, we further extend the $\Omega(\epsilon^{-4})$ lower bound for SGD [3] from single-level optimization to bilevel optimization. Note that existing constructions for bilevel lower bound [13, 29] do not satisfy all our smoothness conditions in Definition 2.2. We demonstrate in Theorem 4.1 that a fully separable construction for upper- and lower-level variables can immediately yield a valid $\Omega(\epsilon^{-4})$ lower bound for the problem class we study, showing that F2SA-$p$ is optimal up to logarithmic factors when $p=\Omega(\log\epsilon^{-1}/\log\log\epsilon^{-1})$ (see Remark 3.5). We summarize our main results, including both the lower and upper bounds, in Table 1 and discuss open problems in the following.
Open problems.
Our upper bounds improve known results for high-order smooth problems, but our result still has a gap between the lower bound for $p={\mathcal{O}}(\log\epsilon^{-1}/\log\log\epsilon^{-1})$. Recently, Kwon et al. [29] obtained some preliminary results towards closing this gap for $p=1$, where they showed an $\Omega(\epsilon^{-6})$ lower bound holds under a more adversarial oracle. But it is still open whether their lower bounds can be extended to standard stochastic oracles as they conjectured. Another open problem is the tightness of the condition number dependency shown in Table 1.
Method Smoothness Reference Complexity
F2SA 1st-order [28] $\tilde{\mathcal{O}}({\rm poly}(\kappa)\epsilon^{-7})$
F2SA 1st-order [29] $\tilde{\mathcal{O}}({\rm poly}(\kappa)\epsilon^{-6})$
F2SA 1st-order [10] $\tilde{\mathcal{O}}(\kappa^{12}\epsilon^{-6})$
F2SA-pp 1st-order Theorem 3.1 $\tilde{\mathcal{O}}(p\kappa^{9+2/p}\epsilon^{-4-2/p})$
Lower Bound $p$th-order in $\bm{y}$} Theorem 4.1 $\Omega(\epsilon^{-4})$
Table 1: The SFO complexity of different methods to find an $\epsilon$-stationary point for $p$th-order smooth first-order bilevel problems with condition number $\kappa$ under standard SGD assumptions.
Notations.
[…]
2 Preliminaries
The goal of bilevel optimization is to minimize the hyper-objective $\varphi({\bm{x}})$, which is in general nonconvex. Since finding a global minimizer of a general nonconvex function requires exponential complexity in the worst case [37, § 1.6], we follow the literature [6, 7] to consider the task of finding an approximate stationary point.
Definition 2.1.
Let $\varphi:{\mathbb{R}}^{d_{x}}\rightarrow{\mathbb{R}}$ be the hyper-objective defined in Eq. (1). We say ${\bm{x}}\in{\mathbb{R}}^{d_{x}}$ is an $\epsilon$-hyper-stationary point if $\|\nabla\varphi({\bm{x}})\|\leq\epsilon$.
Next, we introduce the assumptions used in this paper, which ensure the tractability of the above hyper-stationarity. Compared to [28, 10], we additionally assume the high-order smoothness in lower-level variable $\bm{y}$} to achieve further acceleration.
2.1 Problem Setup
First of all, we follow the standard assumptions on SGD [3] to assume that the stochastic gradient estimators satisfy the following assumption.
Assumption 2.1.
There exists stochastic gradient estimators $F({\bm{x}},{\bm{y}})$ and $G({\bm{x}},{\bm{y}})$ such that
$\mathbb{E}F({\bm{x}},{\bm{y}};\xi)=\nabla f({\bm{x}},{\bm{y}}),\quad\mathbb{E}\|F({\bm{x}},{\bm{y}})-\nabla f({\bm{x}},{\bm{y}})\|^{2}\leq\sigma^{2};$
$\mathbb{E}G({\bm{x}},{\bm{y}};\zeta)=\nabla g({\bm{x}},{\bm{y}}),\quad\mathbb{E}\|G({\bm{x}},{\bm{y}})-\nabla g({\bm{x}},{\bm{y}})\|^{2}\leq\sigma^{2},$
where $\sigma>0$ is the variance of the stochastic gradient estimators. We also partition $F=(F_{x},F_{y})$ and $G=(G_{x},G_{y})$ such that $F_{x},F_{y},G_{x},G_{y}$ are estimators to $\nabla_{x}f,\nabla_{y}f,\nabla_{x}g,\nabla_{y}g$, respectively.
Second, we assume that the hyper-objective $\varphi({\bm{x}})=f({\bm{x}},{\bm{y}}^{*}({\bm{x}}))$ is lower bounded. Otherwise, any algorithm requires infinite time to find a stationary point. Note that the implicit condition $\inf_{{\bm{x}}\in{\mathbb{R}}^{d_{x}}}\varphi({\bm{x}})>$-$\infty$ below can also be easily implied by a more explicit condition on the lower boundedness of upper-level function $\inf_{{\bm{x}}\in{\mathbb{R}}^{d_{x}},{\bm{y}}\in{\mathbb{R}}^{d_{y}}}f({\bm{x}},{\bm{y}})>$-$\infty$.
Assumption 2.2.
The hyper-objective defined in Eq. (1) is lower bounded, and we have
$\varphi({\bm{x}}_{0})-\inf_{{\bm{x}}\in{\mathbb{R}}^{d_{x}}}\varphi({\bm{x}})\leq\Delta,$
where $\Delta>0$ is the initial suboptimality gap and we assume ${\bm{x}}_{0}={\bm{0}}$ without loss of generality.
Third, we assume the lower-level function $g({\bm{x}},{\bm{y}})$ is strongly convex in $\bm{y}$}. It guarantees the uniqueness of ${\bm{y}}^{*}({\bm{x}})$ and the tractability of the bilevel problem. Although not all the problems in applications satisfy the lower-level strong convexity assumption, it is impossible to derive dimension-free upper bounds when the lower-level problem is only convex [9, Theorem 3.2]. Hence, we follow most existing works to consider strongly convex lower-level problems.
Assumption 2.3.
$g({\bm{x}},{\bm{y}})$ is $\mu$-strongly convex in $\bm{y}$}, i.e., for any ${\bm{y}}_{1},{\bm{y}}_{2}\in{\mathbb{R}}^{d_{y}}$, we have
$g({\bm{x}},{\bm{y}}_{2})\geq g({\bm{x}},{\bm{y}}_{1})+\langle\nabla_{y}g({\bm{x}},{\bm{y}}_{1}),{\bm{y}}_{2}-{\bm{y}}_{1}\rangle+\frac{\mu}{2}\|{\bm{y}}_{1}-{\bm{y}}_{2}\|^{2},$
where $\mu>0$ is the strongly convex parameter.
Fourth, we require the following smoothness assumptions following [17]. According to Eq. (2), these conditions are necessary and sufficient to guarantee the Lipschitz continuity of $\nabla\varphi({\bm{x}})$, which further ensure the tractability of an approximate stationary point of the nonconvex hyper-objective $\varphi({\bm{x}})$ [48, 27].
Assumption 2.4.
For the upper-lower function $f$ and lower-level function $g$, we assume that
  1. 1.
    $f({\bm{x}},{\bm{y}})$ is $L_{0}$-Lipschitz in $\bm{y}$}.
  2. 2.
    $\nabla f({\bm{x}},{\bm{y}})$ and $\nabla g({\bm{x}},{\bm{y}})$ are $L_{1}$-Lipschitz jointly in $({\bm{x}},{\bm{y}})$.
  3. 3.
    $\nabla_{xy}^{2}g({\bm{x}},{\bm{y}})$ and $\nabla_{yy}^{2}g({\bm{x}},{\bm{y}})$ are $L_{2}$-Lipschitz jointly in $({\bm{x}},{\bm{y}})$.
We refer to the problem class that jointly satisfies all the above Assumption 2.1, 2.2, 2.3 and 2.4 as first-order smooth bilevel problems, for which [29, 10] showed the F2SA method achieves the $\tilde{\mathcal{O}}(\epsilon^{-6})$ upper complexity bound. In this work, we show an improved bound under the following additional higher-order smoothness assumption on lower-level variable $\bm{y}$}.
Assumption 2.5 (High order smoothness in $y$).
Given $p\in{\mathbb{N}}_{+}$, we assume that
  1. 1.
    $\frac{\partial^{q}}{\partial{\bm{y}}^{q}}\nabla f({\bm{x}},{\bm{y}})$ is $L_{q+1}$-Lipschitz for all $q=1,\cdots,p-1$.
  2. 2.
    $\frac{\partial^{q+1}}{\partial{\bm{y}}^{q+1}}\nabla g({\bm{x}},{\bm{y}})$ is $L_{q+2}$-Lipschitz in $\bm{y}$} for all $q=1,\cdots,p-1$.
We refer to problems jointly satisfying all the above assumptions as $p$th-order smooth bilevel problems, and also formally define their condition numbers as follows.
Definition 2.2 ($p$th-order smooth bilevel problems).
Given $p\in{\mathbb{N}}_{+}$, $\Delta>0$, $L_{0},L_{1},\cdots,L_{p+1}>0$, and $\mu\leq L_{1}$, we use ${\mathcal{F}}^{\text{nc-sc}}(L_{0},\cdots,L_{p+1},\mu,\Delta)$ to denote the set of all bilevel instances satisfying Assumption 2.2, 2.3, 2.4 and 2.5. For this problem class, we define the largest smoothness constant $\bar{L}=\max_{0\leq j\leq p}L_{j}$ and condition number $\kappa=\bar{L}/\mu$.
2.2 Comparison to Previous Works
All our above assumptions align with [10] except for the additional Assumption 2.5. Since we are not the first work to demonstrate that additional assumptions can lead to acceleration in bilevel optimization, we first give a detailed discussion on other assumptions made in related works to see our differences before we show our improved upper bound.
Stochastic Hessian assumption.
Ghadimi and Wang [17], Ji et al. [21] assumes the access to a stochastic Hessian estimator ${\bm{H}}({\bm{x}},{\bm{y}})$ such that
$\mathbb{E}{\bm{H}}({\bm{x}},{\bm{y}})=\nabla^{2}g({\bm{x}},{\bm{y}}),\quad\mathbb{E}\|{\bm{H}}({\bm{x}},{\bm{y}})-\nabla^{2}g({\bm{x}},{\bm{y}})\|\leq\sigma^{2}.$ (5)
Under this assumption, in conjunction with Assumption 2.2, 2.3, and 2.4, Ghadimi and Wang [17] proposed the BSA method and showed that it finds an $\epsilon$-stationary point of $\varphi({\bm{x}})$ with $\tilde{\mathcal{O}}(\epsilon^{-6})$ stochastic gradient oracles and $\tilde{\mathcal{O}}(\epsilon^{-4})$ stochastic HVP oracles. Later, Ji et al. [21] proposed the stocBiO method that requires $\tilde{\mathcal{O}}(\epsilon^{-4})$ stochastic gradient and HVP oracles. Compared to them, we consider the setting where the algorithms only have access to stochastic gradient estimators, and we make no assumptions on the stochastic Hessians.
Mean-squared smoothness assumption.
Besides Assumption 2.1, 2.2, 2.3, 2.4 and the stochastic Hessian assumption (5), Khanduri et al. [25], Yang et al. [46, 47] further assumes that the stochastic estimators to gradients and Hessians are mean-squared smooth:
$\begin{split}\mathbb{E}\|F({\bm{x}},{\bm{y}})-F({\bm{x}}^{\prime},{\bm{y}}^{\prime})\|^{2}\leq\bar{L}_{1}^{2}\|({\bm{x}},{\bm{y}})-({\bm{x}}^{\prime},{\bm{y}}^{\prime})\|^{2},\\ \mathbb{E}\|G({\bm{x}},{\bm{y}})-G({\bm{x}}^{\prime},{\bm{y}}^{\prime})\|^{2}\leq\bar{L}_{1}^{2}\|({\bm{x}},{\bm{y}})-({\bm{x}}^{\prime},{\bm{y}}^{\prime})\|^{2},\\ \mathbb{E}\|{\bm{H}}({\bm{x}},{\bm{y}})-{\bm{H}}({\bm{x}}^{\prime},{\bm{y}}^{\prime})\|^{2}\leq\bar{L}_{2}^{2}\|({\bm{x}},{\bm{y}})-({\bm{x}}^{\prime},{\bm{y}}^{\prime})\|^{2}.\end{split}$ (6)
Under this additional assumption, they proposed faster stochastic methods with upper complexity bound of $\tilde{\mathcal{O}}(\epsilon^{-3})$ via variance reduction [15, 12]. However, variance reduction are typically ineffective in practice [14] since the mean-squared smoothness constants $\bar{L}_{1}$ and $\bar{L}_{2}$ can be arbitrarily worse than the smoothness constants $L_{1}$ and $L_{2}$. In this paper, we only consider the setting without mean-squared smoothness assumptions and study a different acceleration mechanism from variance reduction.
Jointly high-order smoothness assumption.
Huang et al. [20] introduced a second-order smoothness assumption similar to but stronger than Assumption 2.5 when $p=2$. Specifically, they assumed the second-order smoothness jointly in $({\bm{x}},{\bm{y}})$ instead of $\bm{y}$} only:
[…]
The jointly second-order smoothness (7) ensures that the hyper-objective $\varphi({\bm{x}})$ has Lipschitz continuous Hessians, which further allows the application of known techniques in minimizing second-order smooth objectives. Huang et al. [20] applied the technique from [23, 24, 44, 1] to show that an HVP-based method can find a second-order stationary point in $\tilde{\mathcal{O}}(\epsilon^{-2})$ complexity under the deterministic setting, and in $\tilde{\mathcal{O}}(\epsilon^{-4})$ under the stochastic Hessian assumption (5). Yang et al. [45] applied the technique from [31] to accelerate the complexity HVP-based method to $\tilde{\mathcal{O}}({\epsilon}^{-1.75})$ in the deterministic setting. Chen et al. [10] also proposed a fully first-order method to achieve the same $\tilde{\mathcal{O}}({\epsilon}^{-1.75})$ complexity. Compared to these works, our work demonstrates a unique acceleration mechanism in stochastic bilevel optimization that only comes from the high-order smoothness in $\bm{y}$}.
3 The F2SA-$p$ Method
In this section, we introduce our improved method for stochastic first-order bilevel optimization. In Section 3.1, we establish the relationship between fully first-order hyper-gradient approximation and finite difference schemes. In Section 3.2, we design our algorithm inspired by the $p$th-order finite difference method. In Section 3.3, we present the theoretical analysis of our algorithm and show that it achieves a $\tilde{\mathcal{O}}(p\epsilon^{-4-2/p})$ upper bound on SFO calls.
3.1 Hyper-Gradient Approximation via Finite Difference
The core idea of F2SA [28] is to solve the reformulated penalty problem (3) and use the gradient of the penalty function to approximate the true hyper-gradient. In this subsection, we revisit their hyper-gradient approximation and make connections to the finite difference method, which further motivates us to design better algorithms by using better finite difference formulas.
We start from the following relationship of the difference between perturbed lower-level problems and the original problem. We let
$g_{\nu}({\bm{x}},{\bm{y}})$ $:=\nu f({\bm{x}},{\bm{y}})+g({\bm{x}},{\bm{y}}),$
${\bm{y}}_{\nu}^{*}({\bm{x}})$ $:=\arg\min_{{\bm{y}}\in{\mathbb{R}}^{d_{y}}}g_{\nu}({\bm{x}},{\bm{y}}),$
$\ell_{\nu}({\bm{x}})$ $:=\min_{{\bm{y}}\in{\mathbb{R}}^{d_{y}}}g_{\nu}({\bm{x}},{\bm{y}}),$
where $g_{\nu}$ is the perturbed lower-lever problem with ${\bm{y}}_{\nu}^{*}({\bm{x}})$ and $\ell_{\nu}({\bm{x}})$ being its optimal solution and optimal value, respectively. Since the constraint ${\bm{y}}=\arg\min_{{\bm{z}}\in{\mathbb{R}}^{d_{z}}}g({\bm{x}},{\bm{z}})$ is equivalent to requiring $g({\bm{x}},{\bm{y}})\leq\min_{{\bm{z}}\in{\mathbb{R}}^{d_{y}}}g({\bm{x}},{\bm{z}})$, it can be shown [10, Lemma B.3] that $\frac{\partial}{\partial\nu}\ell_{\nu}({\bm{x}})|_{\nu=0}=\varphi({\bm{x}})$ holds under Assumption 2.3 and 2.4. Furthermore, Kwon et al. [28] showed that the partial derivatives with respect to $\bm{x}$} and $\nu$ are commutable, which leads to
$\frac{\partial^{2}}{\partial\nu\partial{\bm{x}}}\ell_{\nu}({\bm{x}})|_{\nu=0}=\frac{\partial^{2}}{\partial{\bm{x}}\partial\nu}\ell_{\nu}({\bm{x}})|_{\nu=0}=\nabla\varphi({\bm{x}}).$ (8)
Let $\nu=1/\lambda$ in Eq. (3). Then the fully first-order hyper-gradient estimator [28, 10] is exactly using forward difference to approximate $\nabla\varphi({\bm{x}})$, that is,
$\frac{\frac{\partial}{\partial{\bm{x}}}\ell_{\nu}({\bm{x}})-\frac{\partial}{\partial{\bm{x}}}\ell_{0}({\bm{x}})}{\nu}\approx\frac{\partial^{2}}{\partial\nu\partial{\bm{x}}}\ell_{\nu}({\bm{x}})|_{\nu=0}=\nabla\varphi({\bm{x}}).$ (9)
However, the forward difference is not the only way to approximate a derivative. Essentially, it falls into a general class of $p$th-order finite difference [4] that can guarantee an ${\mathcal{O}}(\nu^{p})$ approximation error. We restate this known result in the following lemma and provide a self-contained proof in Appendix B for completeness.
Lemma 3.1.
Assume the unitary function $\psi:{\mathbb{R}}\rightarrow{\mathbb{R}}$ has $C$-Lipschitz continuous $p$th-order derivative. If $p$ is even, there exist $p$th-order central difference coefficients $\{\alpha_{j}\}_{j=-p/2}^{p/2}$ such that
$\left|\frac{1}{\nu}\sum_{j=-p/2}^{p/2}\alpha_{j}\psi(j\nu)-\psi^{\prime}(0)\right|={\mathcal{O}}(C\nu^{p}),$
where $\alpha_{0}=0$ and $\alpha_{j}=\alpha_{-j}$ for all $j=1,\cdots,p/2$. If $p$ is odd, there exist $p$th-order forward difference coefficients $\{\beta_{j}\}_{j=0}^{p}$ such that
$\left|\frac{1}{\nu}\sum_{j=-p/2}^{p/2}\beta_{j}\psi(j\nu)-\psi^{\prime}(0)\right|={\mathcal{O}}(C\nu^{p}),$
When $p=1$, we have $\beta_{0}=-1$, $\beta_{1}=1$, and we obtain the forward difference estimator $\psi(\nu)-\psi(0)/\nu$; When $p=2$ we have $\alpha_{-1}=-1/2,\alpha_{1}=1/2$ and we obtain the central difference estimator $(\psi(\nu)-\psi(-\nu))/(2\nu)$. Lemma 3.1 tells us that in general we can always construct a finite difference estimator ${\mathcal{O}}(\nu^{p})$ error with $p$ points with for even $p$ or $p+1$ points for odd $p$ under the given smoothness conditions. Inspired by Lemma 3.1 and Eq. (8) that $\frac{\partial^{2}}{\partial\nu\partial{\bm{x}}}\ell_{\nu}({\bm{x}})|_{\nu=0}=\nabla\varphi({\bm{x}})$, we propose a fully first-order estimator via a linear combination of $\frac{\partial}{\partial{\bm{x}}}\ell_{j\nu}({\bm{x}})$ to achieve ${\mathcal{O}}(\nu^{p})$ approximation error to $\nabla\varphi({\bm{x}})$ given that $\frac{\partial^{p+1}}{\partial\nu^{p}\partial{\bm{x}}}\ell_{\nu}({\bm{x}})$ is Lipschitz continuous in $\nu$. It further leads to Algorithm 1 that will be formally introduced in the next subsection.
Remark 3.1.
A subtlety to use Lemma 3.1 for hyper-gradient estimation is that it only applies to a unitary function while $\frac{\partial}{\partial{\bm{x}}}\ell_{\nu}({\bm{x}})$ is a vector-valued function in $\nu$. However, the approximation error still holds for the whole vector under the Euclidean norm if we apply the lemma on each dimension and note that the finite difference coefficients are the same for all dimensions.
3.2 The Proposed Algorithm
Algorithm 1 F2SA-$p$ $({\bm{x}}_{0},{\bm{y}}_{0})$,   even $p$
[…]
2:for $t=0,1,\cdots,T-1$
3:  Sample random i.i.d indexes $\{(\xi^{y}_{j},\zeta^{y}_{j})\}_{j=-p/2}^{p/2}$ and $\{({\xi_{i}^{x}},{\zeta_{i}^{x}})\}_{i=1}^{S}.$
4:  for $j=-p/2,-p/2+1,\cdots,p/2$
5:   ${\bm{y}}_{t}^{j,0}={\bm{y}}_{t}^{j}$
6:   for $k=0,1,\cdots,K-1$
7:    ${\bm{y}}_{t}^{j,k+1}={\bm{y}}_{t}^{j,k}-\eta_{y}\left(j\nu F_{y}({\bm{x}}_{t},{\bm{y}}_{t}^{j,k};\xi^{y}_{j})+G_{y}({\bm{x}}_{t},{\bm{y}}_{t}^{j,k};\zeta^{y}_{j})\right)$
8:   end for
9:   ${\bm{y}}_{t+1}^{j}={\bm{y}}_{t}^{j,K}$
10:  end for
11:  Let $\{\alpha_{j}\}_{j=-p/2}^{p/2}$ be the $p$th-order central coefficients defined in Lemma 3.1.
12:  $\Phi_{t}=\frac{1}{S}\sum_{i=1}^{S}\sum_{j=-p/2}^{p/2}\alpha_{j}\left(jF_{x}({\bm{x}}_{t},{\bm{y}}_{t+1}^{j};\xi^{x}_{i})+\dfrac{G_{x}({\bm{x}}_{t},{\bm{y}}_{t+1}^{j};\zeta^{x}_{i})}{\nu}\right)$
13:  ${\bm{x}}_{t+1}={\bm{x}}_{t}-\eta_{x}{\Phi_{t}}/{\|\Phi_{t}\|}$
14:end for
Due to space limitations, we only present Algorithm 1 designed for even $p$ in the main text. The algorithm for odd $p$ can be designed similarly, and we defer the concrete algorithm to Appendix D. Our Algorithm 1 follows the double-loop structure of F2SA [10, 29] and changes the hyper-gradient estimator to the one introduced in the previous section. Now, we give a more detailed introduction to the procedures of the two loops of F2SA-$p$.
  1. 1.
    In the outer loop, the algorithm first samples a mini-batch with size $S$ and uses Lemma 3.1 to construct $\Phi_{t}$ via the linear combination of $\frac{\partial}{\partial{\bm{x}}}\ell_{j\nu}({\bm{x}}_{t})$ for $j=-p/2,\cdots,p/2$ every iteration. After obtaining $\Phi_{t}$ as an approximation to $\nabla\varphi({\bm{x}}_{t})$, the algorithm then performs a normalized gradient descent step ${\bm{x}}_{t+1}={\bm{x}}_{t}-\eta_{x}\Phi_{t}/\|\Phi_{t}\|$ with total $T$ iterations.
  2. 2.
    In the inner loop, the algorithm returns an approximation to $\frac{\partial}{\partial{\bm{x}}}\ell_{j\nu}({\bm{x}}_{t})$ for all $j=-p/2,\cdots,p/2$. Note that Danskin’s theorem indicates $\frac{\partial}{\partial{\bm{x}}}\ell_{j\nu}({\bm{x}}_{t})=\frac{\partial}{\partial{\bm{x}}}g_{j\nu}({\bm{x}}_{t},{\bm{y}}_{j\nu}^{*}({\bm{x}}_{t}))$. It suffices to approximate ${\bm{y}}_{j\nu}^{*}({\bm{x}}_{t})$ to sufficient accuracy, which is achieved by taking a $K$-step single-batch SGD subroutine with stepsize $\eta_{y}$ on each function $g_{j\nu}({\bm{x}},\,\cdot\,)$.
Remark 3.2 (Effect of normalized gradient step).
Compared to [10, 28], the only modification we make to the outer loop is to change the gradient step to a normalized gradient step. The normalization can control the change of ${\bm{y}}_{j\nu}^{*}({\bm{x}}_{t})$ and make the analysis of inner loops easier. We believe that all our theoretical guarantees also hold for the standard gradient step via a more involved analysis.
3.3 Complexity Analysis
This section contains the complexity analysis of Algorithm 1. We first derive the following lemma from the high-dimensional Faà di Bruno formula [32].
Lemma 3.2.
Let $\nu\in(0,1/(2\kappa)]$. For any instance in ${\mathcal{F}}^{\text{nc-sc}}(L_{0},\cdots,L_{p+1},\mu,\Delta)$ as per Definition 2.2, $\frac{\partial^{p+1}}{\partial\nu^{p}\partial{\bm{x}}}\ell_{\nu}({\bm{x}})$ is ${\mathcal{O}}(\kappa^{2p+1}\bar{L})$-Lipschitz continuous in $\nu$.
Our result generalizes the prior result for $p=1$ [28] to any $p\in{\mathbb{N}}_{+}$ and also tightens the prior bounds for $p=2$ [10] as we remark in the following.
Remark 3.3 (Tighter bounds for $p=2$).
Note that the variables $\bm{x}$} and $\nu$ play equal roles in our analysis. Therefore, our result in $p=2$ essentially implies that $\frac{\partial^{3}}{\partial\nu\partial{\bm{x}}^{2}}\ell_{\nu}({\bm{x}})$ is ${\mathcal{O}}(\kappa^{5}\bar{L})$-Lipschitz continuous in $\nu$ around zero, which tightens the ${\mathcal{O}}(\kappa^{6}\bar{L})$ bound of Hessian convergence in [10, Lemma 5.1a] and is of independent interest. The main insight is to avoid the direct calculation of $\nabla^{2}\varphi({\bm{x}})=\frac{\partial^{3}}{\partial\nu\partial{\bm{x}}^{2}}\ell_{\nu}({\bm{x}})|_{\nu=0}$ which involves third-order derivatives and makes the analysis more complex, but instead always to analyze it through the limiting point $\lim_{\nu\rightarrow 0_{+}}\frac{\partial^{3}}{\partial\nu\partial{\bm{x}}^{2}}\ell_{\nu}({\bm{x}})$.
Recall Eq. (8) that $\frac{\partial^{2}}{\partial\nu\partial{\bm{x}}}\ell_{\nu}({\bm{x}})|_{\nu=0}=\nabla\varphi({\bm{x}})$. Then Lemma 3.2, in conjunction with Lemma 3.1, indicates that the $p$th-order finite difference used in F2SA-$p$ guarantees an ${\mathcal{O}}(\nu^{p})$-approximation error to $\nabla\varphi({\bm{x}})$, which always improves the ${\mathcal{O}}(\nu)$-error guarantee of F2SA [28, 10] for any $p\geq 2$. This improved error guarantee means that we can set $\nu={\mathcal{O}}(\epsilon^{1/p})$ to obtain an ${\mathcal{O}}(\epsilon)$-accurate hyper-gradient estimator to $\nabla\varphi({\bm{x}})$, which further leads to the following improved complexity of our algorithm.
Theorem 3.1 (Main theorem).
For any instance in ${\mathcal{F}}^{\text{nc-sc}}(L_{0},\cdots,L_{p+1},\mu,\Delta)$ as per Definition 2.2, set the hyper-parameters as
$\begin{split}&\nu\asymp\min\left\{\frac{R}{\kappa},\left(\frac{\epsilon}{\bar{L}\kappa^{2p+1}}\right)^{1/p}\right\},~~\eta_{x}\asymp\frac{\epsilon}{L_{1}\kappa^{3}},~~\eta_{y}\asymp\frac{\nu^{2}\epsilon^{2}}{L_{1}\kappa\sigma^{2}},\\ &\quad\quad S\asymp\frac{\sigma^{2}}{\nu^{2}\epsilon^{2}},~~K\asymp\frac{\kappa^{2}\sigma^{2}}{\nu^{2}\epsilon^{2}}\log\left(\frac{RL_{1}\kappa}{\nu\epsilon}\right),~~T\asymp\frac{\Delta}{\eta_{x}\epsilon},\end{split}$ (10)
where $R=\|{\bm{y}}_{0}-{\bm{y}}^{*}({\bm{x}}_{0})\|$. Run Algorithm 1 if $p$ is even or Algorithm 2 (in Appendix D) if $p$ is odd. Then we can provably find an $\epsilon$-stationary point of $\varphi({\bm{x}})$ with the total SFO calls upper bounded by
$pT(S+K)={\mathcal{O}}\left(\frac{p\Delta L_{1}\bar{L}^{2/p}\sigma^{2}\kappa^{9+2/p}}{\epsilon^{4+2/p}}\log\left(\frac{RL_{1}\bar{L}\kappa}{\epsilon}\right)\right).$
The above theorem shows that the F2SA-$p$ method can achieve the $\tilde{\mathcal{O}}(p\kappa^{9+2/p}\epsilon^{-4-2/p}\log(\nicefrac{{\kappa}}{{\epsilon}}))$ SFO complexity for $p$th-order smooth bilevel problems. In the following, we give several remarks on the complexity in different regions of $p$.
Remark 3.4 (First-order smooth region).
For $p=1$, our upper bound becomes $\tilde{\mathcal{O}}(\kappa^{11}\epsilon^{-6})$, which improves the $\tilde{\mathcal{O}}(\kappa^{12}\epsilon^{-6})$ bound in [10] by a factor of $\kappa$. The improvement comes from a tighter analysis in the lower-level SGD update and a careful parameter setting.
Remark 3.5 (Highly smooth region).
For $p=\Omega(\log(\nicefrac{{\kappa}}{{\epsilon}})/\log\log(\nicefrac{{\kappa}}{{\epsilon}}))$ in Definition 2.2, we can run F2SA-$q$ with $q\asymp\log(\nicefrac{{\kappa}}{{\epsilon}})/\log\log(\nicefrac{{\kappa}}{{\epsilon}})$ and the ${\mathcal{O}}(q\kappa^{9}\epsilon^{-4}(\nicefrac{{\kappa}}{{\epsilon}})^{2/q}\log(\nicefrac{{\kappa}}{{\epsilon}}))$ complexity in Theorem 3.1 simplifies to ${\mathcal{O}}(\kappa^{9}\epsilon^{-4}\log^{3}(\nicefrac{{\kappa}}{{\epsilon}})/\log\log(\nicefrac{{\kappa}}{{\epsilon}}))=\tilde{\mathcal{O}}(\kappa^{9}\epsilon^{-4})$, which matches the best-known complexity for HVP-based methods [21] under stochastic Hessian assumption (5).
In the upcoming section, we will derive an $\Omega(\epsilon^{-4})$ lower bound to prove that the F2SA-$p$ is near-optimal in the above highly smooth region if the condition number $\kappa$ is constant. We leave the study of optimal complexity for non-constant $\kappa$ to future work.
Comparison of results for odd $p$ and even $p$.
Note that by Lemma 3.1 when $p$ is odd, we need to use $p+1$ points to construct the estimator, which means the algorithm needs to solve $p+1$ lower-level problems in each iteration to achieve an ${\mathcal{O}}(\nu^{p})$ error guarantee. In contrast, when $p$ is even, $p$ points are enough since the $p$th-order central difference estimator satisfies that $\alpha_{0}=0$. It suggests that even when $p$ is odd, the algorithm designed for odd $p$ may still be better. For instance, the F2SA-2 may always be a better choice than F2SA since its benefits almost come for free: (1) it still only needs to solve 2 lower-level problems as the F2SA method, which means the per-iteration complexity remains the same. (2) Although the improved complexity of F2SA-2 relies on the second-order smooth condition, without such a condition, its error guarantee in hyper-gradient estimation only degenerates to a first-order one, which means it is at least as good as F2SA.
4 An $\Omega(\epsilon^{-4})$ Lower Bound
In this section, we prove an $\Omega(\epsilon^{-4})$ lower bound for stochastic bilevel optimization via a reduction to single-level optimization. Our lower bound holds for any randomized algorithms A, which consists of a sequence of measurable mappings $\{\texttt{A}_{t}\}_{t=1}^{T}$ that is defined recursively by
$({\bm{x}}_{t+1},{\bm{y}}_{t+1})=\texttt{A}_{t}\left(r,F({\bm{x}}_{0},{\bm{y}}_{0}),G({\bm{x}}_{0},{\bm{y}}_{0})),\cdots,F({\bm{x}}_{t},{\bm{y}}_{t}),G({\bm{x}}_{t},{\bm{y}}_{t}))\right),~~t\in{\mathbb{N}}_{+},$ (11)
where $r$ is a random seed drawn at the beginning to produce the queries, and $F,G$ are the stochastic gradient estimators that satisfy Assumption 2.1. Without loss of generality, we assume that $({\bm{x}}_{0},{\bm{y}}_{0})=({\bm{0}},{\bm{0}})$. Otherwise, we can prove the same lower bound by shifting the functions.
The construction.
We construct a separable bilevel instance such that the upper-level function $f({\bm{x}},{\bm{y}})\equiv f_{\bm{U}}({\bm{x}})$ and its stochastic gradient align with the hard instance in [3], while the lower-level function is the simple quadratic $g({\bm{x}},y)\equiv g(y)=\mu y^{2}/2$ with deterministic gradients. We defer the concrete construction to Appendix E. For this separable bilevel instance, we can show that for any randomized algorithm defined in Eq. (11) that uses oracles $(F_{\bm{U}},G)$, the progress in $\bm{x}$} can be simulated by another randomized algorithm that only uses $F_{\bm{U}}$, meaning that the single-level lower bound [3] also holds.
Theorem 4.1 (Lower bound).
There exist numerical constants $c>0$ such that for all $\Delta>0$, $L_{1},L_{2},\cdots,L_{p+1}>0$ and $\epsilon\leq c\sqrt{L_{1}\Delta}$, there exists a distribution over the function class ${\mathcal{F}}^{\text{nc-sc}}(L_{0},L_{2},\cdots,L_{p+1},\mu,\Delta)$ and the stochastic gradient estimators satisfying Assumption 2.1, such that any randomized algorithm A defined as Eq. (11) can not find an $\epsilon$-stationary point of $\varphi({\bm{x}})=f({\bm{x}},{\bm{y}}^{*}({\bm{x}}))$ in less than $\Omega(\Delta L_{1}\sigma^{2}\epsilon^{-4})$ SFO calls.
The analysis is simple using our fully separable construction $f({\bm{x}},y)=f_{\bm{U}}({\bm{x}})$ and $g(y)=\mu y^{2}/2$. But we are a bit surprised that our straightforward construction is not used in prior works such as [13]. Below, we give a detailed discussion on the constructions in other works.
Comparison to other bilevel lower bounds.
Dagréou et al. [13] proved lower bounds for finite-sum bilevel optimization via a similar reduction to single-level optimization. However, the direct extension of their construction in the fully stochastic setting gives $f({\bm{x}},{\bm{y}})=f_{\bm{U}}({\bm{y}})$ and $g({\bm{x}},{\bm{y}})=({\bm{x}}-{\bm{y}})^{2}$, where the high-order derivatives of $f({\bm{x}},{\bm{y}})$ not ${\mathcal{O}}(1)$-Lipschitz in $\bm{y}$} and thus violates our assumptions. Kwon et al. [29] also proved an $\Omega(\epsilon^{-4})$ lower bound for stochastic bilevel optimization. However, their construction $f({\bm{x}},y)=y$ and $g({\bm{x}},y)=(f_{\bm{U}}({\bm{x}})-y)^{2}$ violate the first-order smoothness of $g({\bm{x}},y)$ in $\bm{x}$} when $y$ is far way from $f_{\bm{U}}({\bm{x}})$. In this work, we use a fully separable construction to avoid all the aforementioned issues in other works.
5 Experiments
In this section, we conduct numerical experiments to verify our theory. We consider the “learn-to-regularize” problem on the “20 Newsgroup” dataset, which is a very standard benchmark in bilevel optimization [19, 21, 33, 10]. In this task, we aim at learning the optimal regularizer for each parameter of a model. We formulate this task into the following bilevel optimization problem:
$\min_{{\bm{x}}\in{\mathbb{R}}^{p}}\ell^{\rm val}({\bm{y}}),\quad{\rm s.t.}\quad{\bm{y}}\in\arg\min_{{\bm{y}}\in{\mathbb{R}}^{q\times p}}\ell^{\rm tr}({\bm{y}})+\|{\bm{W}}_{{\bm{x}}}{\bm{y}}\|^{2},$
where $\bm{x}$} parameterizes the regularization matrix via ${\bm{W}}_{\bm{x}}={\rm diag}(\exp({\bm{x}}))$, $\bm{y}$} parameterizes a linear model that maps $p=130,107$ features to $q=20$ classes, while $\ell^{\rm val}$ and $\ell^{\rm tr}$ denote the validation and training loss, respectively. The whole dataset contains 18,000 samples. We compare our proposed method F2SA-$p$ with both the previous best fully first-order method F2SA and other Hessian-vector-product-based methods stocBiO [21], MRBO and VRBO [46]. We tune $p$ in $\{1,2,\cdots,5\}$ and find that $p=2$ is the optimal choice. One possible reason is that the instance of both $p=1,2$ only requires solving two lower-level problems at each iteration, but the instance of $p\geq 3$ requires solving more than three lower-level problems and may not be concretely efficient. We regard F2SA-2 as an important instantiation of F2SA-$p$ and present its concrete procedure as well as the comparison to F2SA in Appendix F. For all the algorithms, we search the optimal hyperparameters (including $\eta_{x},\eta_{y},\nu$) in a logarithmic scale with base $10$ and present the experiment results in Figure 1, where we also include a line “w/o Reg” that means the baseline without tuning any regularization. It can be observed that: (1) all the hessian-vector-product-based methods are worse than fully first-order methods; (2) the variance reduction technique in VRBO/MRBO is ineffective and may even harm the performance, which also aligns with the findings in [33]; (3) our method F2SA-2 significantly outperforms all the other algorithms. Our preliminary experiment results on the standard benchmark show the potential of F2SA-2 on large-scale bilevel problems.
Figure 1: Performances of different algorithms when learning the optimal regularization.
6 Conclusions and Future Works
This paper proposes a class of fully first-order method F2SA-$p$ that achieves the $\tilde{\mathcal{O}}(p\epsilon^{-4-2/p})$ SFO complexity for $p$th-order smooth bilevel problems. Our result generalized the best-known $\tilde{\mathcal{O}}({\epsilon}^{-6})$ result [29, 10] from $p=1$ to any $p\in{\mathbb{N}}_{+}$. We also complement our result with an $\Omega(\epsilon^{-4})$ lower bound to show that our method is near-optimal when $p=\Omega(\log\epsilon^{-1}/\log\log\epsilon^{-1})$. Nevertheless, a gap still exists when $p$ is small, and we still do not know how to fill it even for the basic setting $p=1$. Another possible direction is to extend our theory to structured nonconvex-nonconvex bilevel problems studied by many recent works [30, 9, 8, 22, 43, 42].
Appendix A Notations for Tensors
We follow the notation of tensors used by Kolda and Bader [26]. For two $p$-way tensors ${\mathcal{X}}\in{\mathbb{R}}^{n_{1}\times n_{2}\times\cdots\times n_{p}}$ and ${\mathcal{Y}}\in{\mathbb{R}}^{n_{1}\times n_{2}\times\cdots\times n_{p}}$, their inner product $z=\langle{\mathcal{X}},{\mathcal{Y}}\rangle$ is defined as
$\langle{\mathcal{X}},{\mathcal{Y}}\rangle=\sum_{i_{1}=1}^{n_{1}}\sum_{i_{2}=1}^{n_{2}}\cdots\sum_{i_{q}=1}^{n_{p}}{\mathcal{X}}_{i_{1},i_{2},\cdots,i_{p}}{\mathcal{Y}}_{i_{1},i_{2},\cdots,i_{p}}.$
For two tensors ${\mathcal{X}}\in{\mathbb{R}}^{n_{1}\times n_{2}\times\cdots n_{p}}$ and ${\mathcal{Y}}\in{\mathbb{R}}^{m_{1}\times m_{2}\cdots\times m_{q}}$, their outer product ${\mathcal{Z}}={\mathcal{X}}\otimes{\mathcal{Y}}$ is a tensor ${\mathcal{Z}}\in{\mathbb{R}}^{n_{1}\times n_{2}\times\cdots\times n_{p}\times m_{1}\times m_{2}\times\cdots\times m_{q}}$ whose elements are defined as
$\left({\mathcal{X}}\otimes{\mathcal{Y}}\right)_{i_{1},i_{2},\cdots,i_{p},j_{1},j_{2},\cdots,j_{q}}={\mathcal{X}}_{i_{1},i_{2},\cdots,i_{p}}{\mathcal{Y}}_{i_{1},i_{2},\cdots,i_{p}}.$
The operator norm of a tensor ${\mathcal{X}}\in{\mathbb{R}}^{n_{1}\times n_{2}\times\cdots\times n_{p}}$ is defined as
$\|{\mathcal{X}}\|=\sup_{\|{\bm{u}}_{i}\|=1,i=1,\cdots,p}\langle{\mathcal{X}},{\bm{u}}_{1}\otimes{\bm{u}}_{2}\otimes\cdots\otimes{\bm{u}}_{p}\rangle.$
Equipped with the notion of norm, we say a mapping ${\mathcal{T}}:{\mathbb{R}}\rightarrow{\mathbb{R}}^{n_{1}\times n_{2}\times\cdots\times n_{p}}$ is $D$-bounded if
$\|{\mathcal{T}}({\bm{x}})\|\leq D,\quad\forall{\bm{x}}\in{\mathbb{R}}.$
We say $\mathcal{T}$} is $C$-Lipschitz continuous if
$\|{\mathcal{T}}({\bm{x}})-{\mathcal{T}}({\bm{y}})\|\leq C\|{\bm{x}}-{\bm{y}}\|,\quad\forall{\bm{x}},{\bm{y}}\in{\mathbb{R}}.$
Appendix B Proof of Lemma 3.1
If $\psi^{(p)}(\nu)$ is $C$-Lipschitz continuous in $\nu$, then by Taylor’s theorem we have
$\psi(\nu)=\psi(0)+\sum_{k=1}^{p}\frac{(j\nu)^{k}}{k!}\psi^{(k)}(0)+{\mathcal{O}}\left(C\nu^{p+1}\right).$ (12)
If $p$ is even, we choose the generalized central difference. If $p$ is odd, we choose the generalized forward difference. Our choices underpin the following proof.
Proof of Lemma 3.1.
We analyze the case when $p$ is even or odd separately.
If $p$ is even.
For the coefficients $\{\alpha_{j}\}_{j=-p/2}^{p/2}$, we set
$\alpha_{j}=\alpha_{-j},\quad\forall j=0,1,\cdots,p/2.$
Then, summing up Eq. (12) with coefficients $\alpha_{j}$ gives
$\frac{1}{\nu}\sum_{j=-p/2}^{j=p/2}\alpha_{j}\psi(j\nu)=\underbrace{2\sum_{j=1}^{p/2}\alpha_{j}\sum_{k=1,3,\cdots}^{p/2-1}\frac{j^{k}\nu^{k-1}}{k!}\psi^{(k)}(0)}_{(*)}+{\mathcal{O}}\left(C\nu^{p}\right).$
To let term (*) be equivalent to $\psi^{\prime}(0)$, we let $\{\alpha_{j}\}_{j=1}^{p/2}$ satisfy the following equations:
$2\sum_{j=1}^{p/2}\alpha_{j}j^{k}={\bm{1}}_{k=1},\quad\forall k=1,3,\cdots,p/2-1,$
which is equivalent to let $\{j\alpha_{j}\}_{j=1}^{p/2}$ satisfy the following linear equation
$\begin{pmatrix}1&1&1&\cdots&1\\ 1^{2}&2^{2}&3^{2}&\cdots&(p/2)^{2}\\ 1^{4}&2^{4}&3^{4}&\cdots&(p/2)^{4}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 1^{p/2-2}&2^{p/2-2}&3^{p/2-2}&\cdots&(p/2)^{p/2-2}\end{pmatrix}\begin{pmatrix}\alpha_{1}\\ 2\alpha_{2}\\ 3\alpha_{3}\\ \vdots\\ (p/2)\alpha_{p/2}\end{pmatrix}=\begin{pmatrix}1/2\\ 0\\ 0\\ \vdots\\ 0\end{pmatrix}$.
Since the coefficient matrix in the above linear equation is a Vandermonde matrix, we know this equation has a unique solution, which gives the value of $\{\alpha_{j}\}_{j=1}^{p/2}$.
If $p$ is odd.
For the coefficients $\{\beta_{j}\}_{j=0}^{p}$, we first let them satisfy the constraint $\sum_{j=0}^{p}\beta_{j}=0$. Then, summing up Eq. (12) with coefficients $\beta_{j}$ gives
$\frac{1}{\nu}\sum_{j=0}^{j=p}\beta_{j}\psi(j\nu)=\underbrace{\sum_{j=0}^{p}\beta_{j}\sum_{k=1}^{p}\frac{j^{k}\nu^{k-1}}{k!}\psi^{(k)}(0)}_{(*)}+{\mathcal{O}}\left(C\nu^{p}\right).$
To let term (*) be equivalent to $\psi^{\prime}(0)$, we let $\{\beta_{j}\}_{j=0}^{p}$ satisfy the following equations:
$\sum_{j=1}^{p}\beta_{j}j^{k}={\bm{1}}_{k=1},\quad\forall k=1,2,\cdots,p,$
which is equivalent to let $\{j\beta\}_{j=1}^{p}$ satisfy the following linear equation
$\begin{pmatrix}1&1&1&\cdots&1\\ 1&2&3&\cdots&p\\ 1^{2}&2^{2}&3^{2}&\cdots&p^{2}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 1^{p-1}&2^{p-1}&3^{p-1}&\cdots&p^{p-1}\end{pmatrix}\begin{pmatrix}\beta_{1}\\ 2\beta_{2}\\ 3\beta_{3}\\ \vdots\\ p\beta_{p}\end{pmatrix}=\begin{pmatrix}1\\ 0\\ 0\\ \vdots\\ 0\end{pmatrix}$.
As before, the coefficient matrix in the above linear equation is also a Vandermonde matrix. Therefore, we know this equation has a unique solution, which gives the value of $\{\beta_{j}\}_{j=1}^{p}$ and the coefficient $\beta_{0}$ can be calculated by $\beta_{0}=1-\sum_{j=1}^{p}\beta_{j}$. ∎
Appendix C Proof of Lemma 3.2
The proof relies on the high-dimensional version of the Faà di Bruno formula. To formally state the result, we define the following notions. For a mapping ${\mathcal{T}}:{\mathbb{R}}^{m}\rightarrow{\mathbb{R}}^{n_{1}\times\cdots\times n_{q}}$, we define its $k$th-order directional derivative evaluated at ${\bm{z}}\in{\mathbb{R}}^{m}$ along the direction $({\bm{u}}_{1},\cdots,{\bm{u}}_{k})$ as
$\nabla^{k}_{{\bm{u}}_{1},\cdots,{\bm{u}}_{k}}{\mathcal{T}}_{\mid{\bm{z}}}=\nabla^{k}{\mathcal{T}}_{\mid{\bm{z}}}({\bm{u}}_{1},\cdots,{\bm{u}}_{k}).$
We let the symmetric products of ${\bm{u}}_{1},\cdots,{\bm{u}}_{k}$ as
${\bm{u}}_{1}\lor{\bm{u}}_{2}\lor\cdots\lor{\bm{u}}_{k}=\frac{1}{k!}\sum_{\pi\in{\rm Perm}(k)}{\bm{u}}_{\pi(1)}\otimes{\bm{u}}_{\pi(2)}\otimes\cdots\otimes{\bm{u}}_{\pi(k)},$
where $\rm Perm(k)$} denotes the set of permutations of $\{1,2,\cdots,k\}$. Also, we define the set of all (unordered) partitions of a set $A$ into $k$ pairwise disjoint non-empty sets as
${\mathcal{P}}(A,k)=\left\{{\bm{P}}=(P_{1},\cdots,P_{k})\subseteq{\mathcal{B}}(A)\mid A=\cup_{j=1}^{k}P_{j};~\emptyset\notin{\bm{P}};~P_{i}\cap P_{j}=\emptyset,\forall i<j\right\},$
where ${\mathcal{B}}(A)$ is the power set of $A$, i.e., the set of all subsets of $A$. We also abbreviate ${\mathcal{P}}(\{1:q\},k)$ as ${\mathcal{P}}(q,k)$. Using the above notions, we have the following result.
Lemma C.1 ([32, Proposition 3.1]).
Let ${\mathcal{T}}_{1}$ and ${\mathcal{T}}_{2}$ be two mappings. If ${\mathcal{T}}_{1}$ and ${\mathcal{T}}_{2}$ are $k$-times differentiable at the point $\bm{z}$} and ${\mathcal{T}}_{1}({\bm{z}})$, respectively, then the composite mapping ${\mathcal{T}}_{2}\circ{\mathcal{T}}_{1}$ is $k$-times differentiable at the point $\bm{z}$} and we have
$\nabla^{q}({\mathcal{T}}_{2}$ $\circ{\mathcal{T}}_{1})_{\mid{\bm{z}}}(\lor_{i=1}^{q}{\bm{u}}_{i})=\sum_{\begin{subarray}{c}1\leq k\leq q,\\ {\bm{P}}\in{\mathcal{P}}(q,k)\end{subarray}}\nabla^{k}{{\mathcal{T}}_{2}}_{\mid{\mathcal{T}}_{1}({\bm{z}})}\left(\nabla^{|P_{1}|}{{\mathcal{T}}_{1}}_{\mid{\bm{z}}}(\lor_{i\in P_{1}}{\bm{u}}_{i}),\cdots\nabla^{|P_{k}|}{{\mathcal{T}}_{1}}_{\mid{\bm{z}}}(\lor_{i\in P_{k}}{\bm{u}}_{i})\right).$
Recall Danskin’s theorem that $\frac{\partial}{\partial{\bm{x}}}\ell_{\nu}({\bm{x}})=\frac{\partial}{\partial{\bm{x}}}g_{\nu}({\bm{x}},{\bm{y}}_{\nu}^{*}({\bm{x}}))$. We can apply Lemma C.1 with ${\mathcal{T}}_{1}={\bm{y}}_{\nu}^{*}({\bm{x}})$ and ${\mathcal{T}}_{1}=\frac{\partial}{\partial{\bm{x}}}g_{\nu}({\bm{x}},{\bm{y}})$ to obtain that
$\frac{\partial^{q+1}}{\partial\nu^{q}\partial{\bm{x}}}\ell_{\nu}({\bm{x}})$ $=\sum_{\begin{subarray}{c}1\leq k\leq q,\\ {\bm{P}}\in{\mathcal{P}}(q,k)\end{subarray}}\frac{\partial^{k+1}}{\partial{\bm{y}}^{k}\partial{\bm{x}}}g_{\nu}({\bm{x}},{\bm{y}}_{\nu}^{*}({\bm{x}}))\left(\frac{\partial^{|P_{1}|}}{\partial\nu^{|P_{1}|}}{\bm{y}}_{\nu}^{*}({\bm{x}}),\cdots,\frac{\partial^{|P_{k}|}}{\partial\nu^{|P_{k}|}}{\bm{y}}_{\nu}^{*}({\bm{x}})\right).$ (13)
Symmetrically, using the first-order optimality condition $\frac{\partial}{\partial{\bm{y}}}g_{\nu}({\bm{x}},{\bm{y}}_{\nu}^{*}({\bm{x}}))=0$ and where the first identity uses the Lemma C.1 with ${\mathcal{T}}_{1}={\bm{y}}_{\nu}^{*}({\bm{x}})$ and ${\mathcal{T}}_{1}=\frac{\partial}{\partial{\bm{y}}}g_{\nu}({\bm{x}},{\bm{y}})$ yields that
$0$ $=\sum_{\begin{subarray}{c}1\leq k\leq q,\\ {\bm{P}}\in{\mathcal{P}}(q,k)\end{subarray}}\frac{\partial^{k+1}}{\partial{\bm{y}}^{k+1}}g_{\nu}({\bm{x}},{\bm{y}}_{\nu}^{*}({\bm{x}}))\left(\frac{\partial^{|P_{1}|}}{\partial\nu^{|P_{1}|}}{\bm{y}}_{\nu}^{*}({\bm{x}}),\cdots,\frac{\partial^{|P_{k}|}}{\partial\nu^{|P_{k}|}}{\bm{y}}_{\nu}^{*}({\bm{x}})\right).$ (14)
Since ${\mathcal{P}}(q,1)$ contains only one element, the above identity implies that
$\begin{split}\frac{\partial^{q}}{\partial\nu^{q}}{\bm{y}}_{\nu}^{*}({\bm{x}})&=-\left(\nabla^{2}_{yy}g_{\nu}({\bm{x}},{\bm{y}}_{\nu}^{*}({\bm{x}}))\right)^{-1}\sum_{\begin{subarray}{c}2\leq k\leq q,\\ {\bm{P}}\in{\mathcal{P}}(q,k)\end{subarray}}{\bm{w}}_{k,{\bm{P}}},\\ {\rm where}~~{\bm{w}}_{k,{\bm{P}}}&=\frac{\partial^{k+1}}{\partial{\bm{y}}^{k+1}}g_{\nu}({\bm{x}},{\bm{y}}_{\nu}^{*}({\bm{x}}))\left(\frac{\partial^{|P_{1}|}}{\partial\nu^{|P_{1}|}}{\bm{y}}_{\nu}^{*}({\bm{x}}),\cdots,\frac{\partial^{|P_{k}|}}{\partial\nu^{|P_{k}|}}{\bm{y}}_{\nu}^{*}({\bm{x}})\right).\end{split}$ (15)
Based on Eq. (15), we can prove by induction that $\frac{\partial^{q}}{\partial\nu^{q}}{\bm{y}}_{\nu}^{*}({\bm{x}})$ is ${\mathcal{O}}(\kappa^{2q+1})$-Lipschitz continuous in $\nu$ for all $q=0,\cdots,p$. The induction base for $q=0,1$ is already proved by Chen et al. [10].
Lemma C.2 (Chen et al. [10, Lemma B.2 and B.5]).
Let $\nu\in(0,1/(2\kappa)]$. Under Assumption 2.3 and 2.4, ${\bm{y}}_{\nu}^{*}({\bm{x}})$ and $\frac{\partial}{\partial\nu}{\bm{y}}_{\nu}^{*}({\bm{x}})$ is ${\mathcal{O}}(\kappa)$- and ${\mathcal{O}}(\kappa^{3})$-Lipschitz continuous in $\nu$, respectively.
Since Eq. (15) also involves $(\nabla_{yy}^{2}g_{\nu}({\bm{x}},{\bm{y}}_{\nu}^{*}({\bm{x}})))^{-1}$, we also need the following lemma that gives its boundedness and Lipschitz continuity constants.
Lemma C.3 (Chen et al. [10, Lemma B.1 and Eq. 18]).
Let $\nu\in(0,1/(2\kappa)]$. Under Assumption 2.3 and 2.4, $(\nabla_{yy}g_{\nu}({\bm{x}},{\bm{y}}_{\nu}^{*}({\bm{x}})))^{-1}$ is $2/\mu$-bounded and ${\mathcal{O}}(\kappa^{2}/\mu)$-Lipschitz continuous in $\nu$.
In the remaining proofs, we will use Eq. (15) prove by induction that $\frac{\partial^{q}}{\partial\nu^{q}}{\bm{y}}_{\nu}^{*}({\bm{x}})$ is ${\mathcal{O}}(\kappa^{2q+1})$-Lipschitz continuous in $\nu$, then we can easily use Eq. (13) to show that $\frac{\partial^{q+1}}{\partial\nu^{q}\partial{\bm{x}}}\ell_{\nu}({\bm{x}})$ is ${\mathcal{O}}(\kappa^{2q+1}\bar{L})$-Lipschitz continuous in $\nu$ for all $q=0,\cdots,p$. Note that the computational graph of either $\frac{\partial^{q}}{\partial\nu^{q}}{\bm{y}}_{\nu}^{*}({\bm{x}}))$ or $\frac{\partial^{q+1}}{\partial\nu^{q}\partial{\bm{x}}}\ell_{\nu}({\bm{x}})$ in Eq. (13) or (15) defines a tree, where the root is output, the leaves are inputs, and the other nodes are the intermediate results in the computation. We can analyze the Lipschitz continuities of all the nodes from bottom to top using the following lemma.
Lemma C.4 (Luo et al. [35, Lemma 12]).
Let ${\mathcal{T}}_{1}$ and ${\mathcal{T}}_{2}$ be two tensor-to-tensor mappings. If ${\mathcal{T}}_{1}$ is $D_{1}$-bounded and $C_{1}$-Lipschitz continuous, ${\mathcal{T}}_{2}$ is $D_{2}$-bounded and $C_{2}$-Lipschitz continuous, then the product mapping ${\mathcal{T}}_{1}\times{\mathcal{T}}_{2}$ is $D_{1}D_{2}$-bounded and $(C_{1}D_{2}+C_{2}D_{1})$-Lipschitz continuous.
Proof of Lemma 3.2.
Now, we formally begin to prove by induction that $\frac{\partial^{q}}{\partial\nu^{q}}{\bm{y}}_{\nu}^{*}({\bm{x}})$ is ${\mathcal{O}}(\kappa^{2q+1})$-Lipschitz continuous in $\nu$ for all $q=0,\cdots,p$. Recall that the induction base follows Lemma C.2. In the following, we use the induction hypothesis that $\frac{\partial^{k}}{\partial\nu^{k}}{\bm{y}}_{\nu}^{*}({\bm{x}}))$ is ${\mathcal{O}}(\kappa^{2k+1})$-Lipschitz continuous in $\nu$ for all $k=0,\cdots,q-1$ to prove that $\frac{\partial^{q}}{\partial\nu^{q}}{\bm{y}}_{\nu}^{*}({\bm{x}}))$ is ${\mathcal{O}}(\kappa^{2q+1})$-Lipschitz continuous in $\nu$. We know that $\frac{\partial^{k+1}}{\partial{\bm{y}}^{k+1}}g_{\nu}({\bm{x}},{\bm{y}}_{\nu}^{*}({\bm{x}}))$ is ${\mathcal{O}}(\bar{L})$-bounded and ${\mathcal{O}}(\kappa\bar{L})$-Lipschitz continuous in $\nu$. Therefore, we can use Lemma C.4 to conclude that each ${\bm{w}}_{k,{\bm{P}}}$ is ${\mathcal{O}}(\kappa^{\sum_{j=1}^{k}(2|P_{j}|-1)}\bar{L})={\mathcal{O}}(\kappa^{2q-k}\bar{L})$-bounded and ${\mathcal{O}}(\bar{L}\cdot\kappa^{2q-k+2}+\kappa\bar{L}\cdot\kappa^{2q-k})={\mathcal{O}}(\kappa^{2q-k+2}\bar{L})$-Lipschitz continuous in $\nu$. It further implies that the summation ${\bm{w}}:=\sum_{\begin{subarray}{c}2\leq k\leq q,{\bm{P}}\in{\mathcal{P}}(q,k)\end{subarray}}{\bm{w}}_{k,{\bm{P}}}$ is ${\mathcal{O}}(\kappa^{2q-2}\bar{L})$-bounded and ${\mathcal{O}}(\kappa^{2q}\bar{L})$-Lipschitz continuous in $\nu$. Then, we can recall Lemma C.3 that $(\nabla_{yy}g_{\nu}({\bm{x}},{\bm{y}}_{\nu}^{*}({\bm{x}})))^{-1}$ is $2/\mu$-bounded and ${\mathcal{O}}(\kappa^{2}/\mu)$-Lipschitz continuous in $\nu$, and use Eq. (15) to finish the induction that $\frac{\partial^{q}}{\partial\nu^{q}}{\bm{y}}_{\nu}^{*}({\bm{x}})=-\left(\nabla^{2}_{yy}g_{\nu}({\bm{x}},{\bm{y}}_{\nu}^{*}({\bm{x}}))\right)^{-1}{\bm{w}}$ is ${\mathcal{O}}(\kappa^{2q+1})$-Lipschitz continuous in $\nu$ for all $q=0,\cdots,p$. Finally, by analogy with the similarity of Eq. (13) and (15), we can follow the same analysis to show that $\frac{\partial^{q+1}}{\partial\nu^{q}\partial{\bm{x}}}\ell_{\nu}({\bm{x}})$ is ${\mathcal{O}}(\kappa^{2q+1}\bar{L})$-Lipschitz continuous in $\nu$ for all $0,\cdots,p$. ∎
Appendix D Proof of Theorem 3.1
Algorithm 2 F2SA-$p$ $({\bm{x}}_{0},{\bm{y}}_{0})$,   odd $p$
[…]
2:for $t=0,1,\cdots,T-1$
3:  Sample random i.i.d indexes $\{(\xi^{y}_{j},\zeta^{y}_{j})\}_{j=0}^{p}$ and $\{({\xi_{i}^{x}},{\zeta_{i}^{x}})\}_{i=1}^{S}.$
4:  for $j=0,\cdots,p$
5:   ${\bm{y}}_{t}^{j,0}={\bm{y}}_{t}^{j}$
6:   for $k=0,1,\cdots,K-1$
7:    ${\bm{y}}_{t}^{j,k+1}={\bm{y}}_{t}^{j,k}-\eta_{y}\left(j\nu F_{y}({\bm{x}}_{t},{\bm{y}}_{t}^{j,k};\xi_{j}^{y})+G_{y}({\bm{x}}_{t},{\bm{y}}_{t}^{j,k};\zeta_{j}^{y})\right)$
8:   end for
9:   ${\bm{y}}_{t+1}^{j}={\bm{y}}_{t}^{j,K}$
10:  end for
11:  Let $\{\beta_{j}\}_{j=0}^{p}$ be the $p$th-order forward difference coefficients defined in Lemma 3.1.
12:  $\Phi_{t}=\frac{1}{S}\sum_{i=1}^{S}\sum_{j=0}^{p}\beta_{j}\left(jF_{x}({\bm{x}}_{t},{\bm{y}}_{t+1}^{j};\xi^{x}_{i})+\dfrac{G_{x}({\bm{x}}_{t},{\bm{y}}_{t+1}^{j};\zeta^{x}_{i})}{\nu}\right)$
13:  ${\bm{x}}_{t+1}={\bm{x}}_{t}-\eta_{x}{\Phi_{t}}/{\|\Phi_{t}\|}$
14:end for
In the main text, we only present the algorithm when $p$ is even. The algorithm when $p$ is odd follows a similar design, which is presented in Algorithm 2 for completeness. Our algorithms consist of a double loop, where the outer loop performs normalized SGD (NSGD) and the inner loop performs SGD. Before we give the formal proof, we first recall the convergence result for (N)SGD.
Lemma D.1 (Cutkosky and Mehta [11, Lemma 2]).
Consider the NSGD update ${\bm{x}}_{t+1}={\bm{x}}_{t}-\eta F_{t}/\|F_{t}\|$ to optimize a function $f:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}$ with $L$-Lipschitz continuous gradients. We have
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\|\nabla f({\bm{x}}_{t}))\|\leq\frac{3(f({\bm{x}}_{0})-\inf_{{\bm{x}}\in{\mathbb{R}}^{d}}f({\bm{x}}))}{\eta T}+\frac{3L\eta}{2}+\frac{8}{T}\sum_{t=0}^{T-1}\mathbb{E}\|F_{t}-\nabla f({\bm{x}}_{t})\|.$
Lemma D.2 (Kwon et al. [29, Lemma C.1]).
Consider the SGD update ${\bm{x}}_{t+1}={\bm{x}}_{t}-\eta F_{t}/\|F_{t}\|$ to optimize a $\mu$-strongly convex function $f:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}$ with $L$-Lipschitz continuous gradients. Let ${\bm{x}}^{*}=\arg\min_{{\bm{x}}\in{\mathbb{R}}^{d}}f({\bm{x}})$ be the unique minimizer to $f$. Suppose $F_{t}$ is an unbiased estimator to $\nabla f({\bm{x}}_{t})$ with variance bounded by $\sigma^{2}$. Setting $\eta<2/(\mu+L)$, we have
$\mathbb{E}\|{\bm{x}}_{t}-{\bm{x}}^{*}\|^{2}\leq(1-\mu\eta)^{t}\|{\bm{x}}_{0}-{\bm{x}}^{*}\|^{2}+\frac{\eta\sigma^{2}}{\mu}.$
The following two lemmas are also useful in the analysis.
Lemma D.3 (Chen et al. [10, Lemma 4.1]).
Under Assumption 2.3, and 2.4, the hyper-objective $\varphi({\bm{x}})=f({\bm{x}},{\bm{y}}^{*}({\bm{x}}))$ is differentiable and has $L_{\varphi}={\mathcal{O}}(\bar{L}\kappa^{3})$-Lipschitz continuous gradients.
Lemma D.4 (Chen et al. [10, Lemma B.6]).
Let $\nu\in(-1/\kappa,1/\kappa)$. Under Assumption 2.3, and 2.4, the optimal (perturbed) lower-level solution mapping ${\bm{y}}_{\nu}^{*}({\bm{x}})=\arg\min_{{\bm{y}}\in{\mathbb{R}}^{d_{y}}}\ell_{v}({\bm{x}},{\bm{y}})$ is $4\kappa$-Lipschitz continuous in $\bm{x}$}.
Now we are ready to prove Theorem 3.1.
Proof of Theorem 3.1.
We separately consider the complexity for the outer loop and the inner loop.
Outer Loop.
According to Lemma D.3, the hyper-objective $\varphi({\bm{x}})$ has $L_{\varphi}={\mathcal{O}}(\bar{L}\kappa^{3})$-Lipschitz continuous gradients. If we can guarantee the condition
$\mathbb{E}\|\Phi_{t}-\nabla\varphi({\bm{x}}_{t})\|\leq\frac{\epsilon}{32},\quad t=0,\cdots,T-1,$ (16)
then we can further set $\eta_{x}=\nicefrac{{\epsilon}}{{6L_{\varphi}}}$ and apply Lemma D.1 to conclude that the algorithm can provably find an $\epsilon$-stationary point of $\varphi({\bm{x}})$ in $T=\left\lceil\nicefrac{{6\Delta}}{{\epsilon\eta_{x}}}\right\rceil={\mathcal{O}}(\Delta L_{1}\kappa^{3}\epsilon^{-2})$ outer iterations.
Inner Loop.
From the above analysis, the remaining goal is to show that the inner loop always returns $\Phi_{t}$ satisfying Eq. (16), which requires $\mathbb{E}\|\Phi_{t}-\nabla\varphi({\bm{x}}_{t})\|={\mathcal{O}}(\epsilon)$ for all $t=0,\cdots,T-1$. Note that the setting of mini-batch size $S=\Omega\left(\nicefrac{{\sigma^{2}}}{{\nu^{2}\epsilon^{2}}}\right)$ ensures that
$\begin{cases}\mathbb{E}\left\|\Phi_{t}-\sum_{j=-p/2}^{p/2}\alpha_{j}\left(j\nabla_{x}f({\bm{x}}_{t},{\bm{y}}_{t+1}^{j})+\dfrac{\nabla_{x}g({\bm{x}}_{t},{\bm{y}}_{t+1}^{j})}{\nu}\right)\right\|={\mathcal{O}}(\epsilon),&\quad p\text{ is even};\\[14.22636pt] \mathbb{E}\left\|\Phi_{t}-\sum_{j=0}^{p}\beta_{j}\left(j\nabla_{x}f({\bm{x}}_{t},{\bm{y}}_{t+1}^{j})+\dfrac{\nabla_{x}g({\bm{x}}_{t},{\bm{y}}_{t+1}^{j})}{\nu}\right)\right\|={\mathcal{O}}(\epsilon),&\quad p\text{ is odd}.\end{cases}$
By Lemma 3.2 and Lemma 3.1, setting $\nu={\mathcal{O}}((\nicefrac{{\epsilon}}{{\bar{L}\kappa^{2p+1}}})^{1/p})$ can ensure that
$\begin{cases}\left\|\nabla\varphi({\bm{x}}_{t})-\sum_{j=-p/2}^{p/2}\alpha_{j}\left(j\nabla_{x}f({\bm{x}}_{t},{\bm{y}}_{j\nu}^{*}({\bm{x}}_{t}))+\dfrac{\nabla_{x}g({\bm{x}}_{t},{\bm{y}}_{j\nu}^{*}({\bm{x}}_{t}))}{\nu}\right)\right\|={\mathcal{O}}(\epsilon),&\quad p\text{ is even};\\[14.22636pt] \left\|\nabla\varphi({\bm{x}}_{t})-\sum_{j=0}^{p}\beta_{j}\left(j\nabla_{x}f({\bm{x}}_{t},{\bm{y}}_{j\nu}^{*}({\bm{x}}_{t}))+\dfrac{\nabla_{x}g({\bm{x}}_{t},{\bm{y}}_{j\nu}^{*}({\bm{x}}_{t}))}{\nu}\right)\right\|={\mathcal{O}}(\epsilon),&\quad p\text{ is odd}.\end{cases}$
Therefore, a sufficient condition of $\mathbb{E}\|\Phi_{t}-\nabla\varphi({\bm{x}}_{t})\|={\mathcal{O}}(\epsilon)$ is
$\begin{split}\begin{cases}\|{\bm{y}}_{t+1}^{j}-{\bm{y}}_{j\nu}^{*}({\bm{x}}_{t})\|={\mathcal{O}}(\nu\epsilon/L_{1}),~~\forall j=-p/2,\cdots,p/2,&\quad p\text{ is even};\\[5.69054pt] \|{\bm{y}}_{t+1}^{j}-{\bm{y}}_{j\nu}^{*}({\bm{x}}_{t})\|={\mathcal{O}}(\nu\epsilon/L_{1}),~~\forall j=0,\cdots,p,&\quad p\text{ is odd}.\end{cases}\end{split}$ (17)
Our next goal is to show that our parameter setting fulfills Eq. (17). Note that for $\nu={\mathcal{O}}(1/\kappa)$, the (perturbed) lower-level problem $g_{j\nu}({\bm{x}},{\bm{y}})$ is $\Omega(\mu)$-strongly convex in $\bm{y}$} and has ${\mathcal{O}}(L_{1})$-Lipschitz continuous gradients jointly in $({\bm{x}},{\bm{y}})$. Therefore, if we set $\eta_{y}\lesssim 1/L_{1}$, then we can apply Lemma D.2 on the lower-level problem $g_{j\nu}({\bm{x}},{\bm{y}})$ to conclude that for ant $j$, we have
$\mathbb{E}\|{\bm{y}}_{t+1}-{\bm{y}}_{j\nu}^{*}({\bm{x}}_{t})\|^{2}\leq(1-\mu\eta_{y})^{K}\|{\bm{y}}_{t}-{\bm{y}}_{j\nu}^{*}({\bm{x}}_{t})\|^{2}+{\mathcal{O}}(\eta_{y}\sigma^{2}/\mu).$
Comparing it with Eq. (17), we can set $\eta_{y}={\mathcal{O}}(\nicefrac{{\nu^{2}\epsilon^{2}}}{{L_{1}\kappa\sigma^{2}}})$ to ensure that for ant $j$, we have
$\mathbb{E}\|{\bm{y}}_{t+1}-{\bm{y}}_{j\nu}^{*}({\bm{x}}_{t})\|\leq(1-\mu\eta_{y})^{K}\|{\bm{y}}_{t}-{\bm{y}}^{*}_{j\nu}({\bm{x}}_{t})\|+{\mathcal{O}}(\nu\epsilon/L_{1}).$
Further, we can use Lemma D.4 and the triangle inequality to obtain that for ant $j$, we have
$\mathbb{E}\|{\bm{y}}_{t+1}-{\bm{y}}_{j\nu}^{*}({\bm{x}}_{t})\|\leq(1-\mu\eta_{y})^{K}(\|{\bm{y}}_{t}-{\bm{y}}_{j\nu}^{*}({\bm{x}}_{t-1})\|+4\kappa\|{\bm{x}}_{t}-{\bm{x}}_{t-1}\|)+{\mathcal{O}}(\nu\epsilon/L_{1}).$ (18)
The recursion (18) implies our setting of $K$ can ensure that Eq. (17) holds for all $t=0,\cdots,T-1$. We give an induction-based proof. To let the induction base holds for $t=1$, it suffices to set $K=\Omega(\nicefrac{{\log(\nicefrac{{RL_{1}}}{{\nu\epsilon}})}}{{\mu\eta_{y}}})=\Omega(\nicefrac{{\log(\nicefrac{{RL_{1}}}{{\nu\epsilon}})\kappa^{2}\sigma^{2}}}{{\nu^{2}\epsilon^{2}}})$, where $\|{\bm{y}}_{j\nu}^{*}({\bm{x}}_{0})-{\bm{y}}^{*}({\bm{x}}_{0})\|^{2}={\mathcal{O}}(R)$ is due to the setting of $\nu={\mathcal{O}}(\nicefrac{{R}}{{\kappa}})$ and the fact that ${\bm{y}}_{\nu}^{*}({\bm{x}})$ is $\kappa$-Lipschitz in $\nu$ by Lemma C.2. Next, assume that we have already guaranteed Eq. (17) holds for iteration $t$, we prove that our setting of $K$ implies Eq. (17) holds for iteration $t+1$. Note that the NSGD update in $\bm{x}$} means that $\|{\bm{x}}_{t}-{\bm{x}}_{t-1}\|=\eta_{x}={\mathcal{O}}(\nicefrac{{\epsilon}}{{6L_{1}\kappa^{3}}})$. Therefore, Eq. (18) in conjunction with the induction hypothesis indicates that
$\mathbb{E}\|{\bm{y}}_{t+1}-{\bm{y}}_{j\nu}^{*}({\bm{x}}_{t})\|\lesssim(1-\mu\eta_{y})^{K}\left(\frac{\nu\epsilon}{L_{1}}+\frac{\epsilon}{L_{1}\kappa^{2}}\right)+\frac{\nu\epsilon}{L_{1}}.$
Therefore, we know that to let Eq. (17) holds for iteration $t+1$, it suffices to let $K=\Omega(\nicefrac{{\log(\nicefrac{{1}}{{\nu\kappa^{2}}})}}{{\mu\eta_{y}}})=\Omega(\nicefrac{{\log(\nicefrac{{1}}{{\nu\kappa^{2}}})\kappa^{2}\sigma^{2}}}{{\nu^{2}\epsilon^{2}}})$. This finishes the induction.
Total Complexity.
According to the above analysis, we set $\nu\asymp(\nicefrac{{\epsilon}}{{\bar{L}\kappa^{2p+1}}})^{1/p}$, $S\asymp\nicefrac{{\sigma^{2}}}{{\nu^{2}\epsilon^{2}}}$, $T\asymp\Delta L_{1}\kappa^{3}\epsilon^{-2}$, and $K\asymp\nicefrac{{\log(\nicefrac{{RL_{1}\kappa}}{{\nu\epsilon}})\kappa^{2}\sigma^{2}}}{{\nu^{2}\epsilon^{2}}}$ to ensure that the algorithm provably find an $\epsilon$-stationary point of $\varphi({\bm{x}})$. Since $S\lesssim K$, the total complexity of the algorithm is
$\quad pT(S+K)={\mathcal{O}}(pTK)={\mathcal{O}}\left(p\cdot\frac{\Delta L_{1}\kappa^{3}}{\epsilon^{2}}\cdot\frac{\kappa^{2}\sigma^{2}}{\nu^{2}\epsilon^{2}}\log\left(\frac{RL_{1}\kappa}{\nu\epsilon}\right)\right)$
$={\mathcal{O}}\left(\frac{p\Delta L_{1}\bar{L}^{2/p}\sigma^{2}\kappa^{9+2/p}}{\epsilon^{4+2/p}}\log\left(\frac{RL_{1}\kappa}{\nu\epsilon}\right)\right).$
Appendix E Proof of Theorem 4.1
We prove our lower bound for stochastic nonconvex-strongly-convex bilevel optimization via a reduction to the lower bound for stochastic single-level nonconvex optimization [3]. To state their lower bound, we first need to introduce the function class, oracle class, algorithm class, and the complexity measures.
Definition E.1.
Given any $\Delta>0$ and $L_{1}>0$, we use ${\mathcal{F}}^{\rm nc}(L_{1},\Delta)$ to denote the set of all smooth functions $f:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}$ that satisfies
  1. 1.
    $f({\bm{0}})-\inf_{{\bm{x}}\in{\mathbb{R}}^{d}}f({\bm{x}})\leq\Delta$;
  2. 2.
    $\nabla f({\bm{x}})$ is $L_{1}$-Lipschitz continuous.
Definition E.2.
Given a function ${\mathbb{R}}^{d}\rightarrow{\mathbb{R}}$, we use ${\mathbb{O}}(\sigma^{2})$ to denote the set of all stochastic first-order oracles that return an unbiased stochastic estimator to $\nabla f$ with variance bounded by $\sigma^{2}$.
Definition E.3.
Let $f:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}$ be a differentiable function and $F:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}$ be the stochastic estimator to $\nabla f$. A randomized first-order algorithm A consists of a distribution ${\mathcal{P}}_{r}$ over a measurable set $\mathcal{R}$} and a sequence of measurable mappings $\{\texttt{A}_{t}\}_{t\in{\mathbb{N}}}$ such that
${\bm{x}}_{t+1}=\texttt{A}_{t}(r,F({\bm{x}}_{0}),\cdots,F({\bm{x}}_{t})),\quad t\in{\mathbb{N}}_{+},$
where $r\sim{\mathcal{P}}_{r}$ is drawn a single time at the beginning of the protocol. We let ${\mathcal{A}}_{\rm rand}$ to denote the class of all the algorithms that follow the above protocol.
Definition E.4.
We define distributional complexity of ${\mathcal{A}}_{\rm rand}$ to find an $\epsilon$-stationary point of the functions in ${\mathcal{F}}^{\rm nc}(L_{1},\Delta)$ with oracle ${\mathbb{O}}(\sigma^{2})$ as
${\rm Compl}_{\epsilon}(L_{1},\Delta,\sigma^{2})=\sup_{\texttt{O}\in{\mathbb{O}}(\sigma^{2})}\sup_{{\mathcal{P}}_{f}\in{\mathcal{P}}[{\mathcal{F}}(\Delta,f)]}\inf_{\texttt{A}\in{\mathcal{A}}_{\rm rand}}\inf\{t\in{\mathbb{N}}\mid\mathbb{E}\|\nabla f({\bm{x}}_{t})\|\leq\epsilon\},$
where the expectation is taken over the sampling of $f$ from ${\mathcal{P}}_{f}$, the randomness in the oracle O, and the randomness in the algorithm A, $\{{\bm{x}}_{t}\}_{t\in{\mathbb{N}}}$ is the sequence generated by A running on $f$ with oracle O, and ${\mathcal{P}}[{\mathcal{F}}^{\rm nc}(L_{1},\Delta)]$ denotes the set of all distributions over ${\mathcal{F}}^{\rm nc}(L_{1},\Delta)$.
All the above definitions are merely restatements of [3, Section 2]. Although Definition E.4 uses the definition of distributional complexity, by Yao’s minimax principle is also a lower bound for the worst-case complexity. Now, we recall the construction in [3] for proving the $\Omega(\epsilon^{-4})$ lower bound. Formally, we define the randomized function
$f_{\bm{U}}({\bm{x}})=\frac{L_{1}\beta^{2}}{\bar{L}_{1}}f^{\rm nc}(\rho({\bm{U}}^{\top}{\bm{x}}/\beta))+\frac{L_{1}\lambda}{2\bar{L}_{1}}\|{\bm{x}}\|^{2},$ (19)
where $\bar{L}_{1}=155$, $\beta=4\bar{L}_{1}\epsilon/L_{1}$, $\rho:{\mathbb{R}}^{T}\rightarrow{\mathbb{R}}^{T}$ is $\rho({\bm{x}})={\bm{x}}\big{/}\sqrt{1+\|{\bm{x}}\|^{2}/R^{2}}$, $R=230\sqrt{T}$, $\lambda=1/5$, and $f_{T}:{\mathbb{R}}^{T}\rightarrow{\mathbb{R}}$ is the nonconvex hard instance introduced by Carmon et al. [6]:
$f^{\rm nc}(x):=-\Psi(1)\Psi(x_{1})+\sum_{i=2}^{T}[\Phi(-x_{i-1})\Phi(-x_{i})-\Psi(x_{i-1})\Phi(x_{i})].$
In the above, the component functions $\Psi,\Phi:{\mathbb{R}}\rightarrow{\mathbb{R}}$ are defined as
$\Psi(t)=\begin{cases}0,&t\leq 1/2,\\ \exp(1-1/(2t-1)^{2}),&t<1/2\end{cases}\quad\text{and}\quad\Phi(t)=\sqrt{\rm e}\int_{-\infty}^{t}\exp(-t^{2}/2){\rm d}t.$
For the hard instance in Eq. (19), Arjevani et al. [3] further defined the stochastic gradient estimator $F_{\bm{U}}$ as
$F_{\bm{U}}({\bm{x}})=\frac{L_{1}}{\bar{L}_{1}}\left(\beta(\nabla\rho({\bm{x}}))^{\top}{\bm{U}}F_{T}({\bm{U}}^{\top}\rho({\bm{x}}))+\lambda{\bm{x}}\right).$ (20)
In the above, $F_{T}:{\mathbb{R}}^{T}\rightarrow{\mathbb{R}}^{T}$ is the stochastic gradient estimator of $\nabla f^{\rm nc}$ defined by
$[F_{T}({\bm{x}})]_{i}=\nabla_{i}f^{\rm nc}({\bm{x}})\left(1+{\bm{1}}_{i>{\rm prog}_{1/4}(x)}(\xi/\gamma-1)\right),\quad\xi\sim{\rm Bernoulli}(\gamma),$
where ${\rm prog}_{\alpha}(x)=\max\{i\geq 0\mid|x_{i}|>\alpha\}$ and $\gamma=\min\{(46\epsilon)^{2}/\sigma^{2},1\}$. For the above construction, Arjevani et al. [3] showed the following lower bound.
Theorem E.1 ([3, Theorem 3]).
There exist numerical constants $c,c^{\prime}>0$ such that for all $\Delta>0$, $L_{1}>0$ and $\epsilon\leq c\sqrt{L_{1}\Delta}$, the construction of function $f_{\bm{U}}:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}$ and stochastic first-order oracle $F_{\bm{U}}:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}$ in Eq. (19) and (20) together give a distribution over the function class ${\mathcal{F}}^{\rm nc}(L_{1},\Delta)$ and a stochastic first-order oracle $\texttt{O}\in{\mathbb{O}}(\sigma^{2})$ such that
${\rm Compl}_{\epsilon}(L_{1},\Delta,\sigma^{2})\geq c^{\prime}\Delta L_{1}\sigma^{2}\epsilon^{-4}.$
Proof of Theorem 4.1.
For any randomized algorithm A defined as Eq. (11) running it on our hard instance, we show that it can be simulated by another randomized algorithm running on the variable $\bm{x}$} such that Theorem E.1 can be applied. Since $G(y)=\mu y$ is a deterministic mapping we know that any randomized algorithm A induces a sequence of measurable mappings $\{\texttt{A}_{t}^{\prime}\}_{t\in{\mathbb{N}}}$ such that
$({\bm{x}}_{t},y_{t})$ $=\texttt{A}_{t}^{\prime}(\xi,F({\bm{x}}_{0}),\cdots,F({\bm{x}}_{t-1}),y_{0},\cdots,y_{t-1}).$
Expanding the recursion for $y_{t}$ shows that the above equation induces another sequence of measurable mappings $\{\texttt{A}_{t}^{\prime\prime}\}_{t\in{\mathbb{N}}}$ such that
$({\bm{x}}_{t},y_{t})=\texttt{A}_{t}^{\prime\prime}(\xi,F({\bm{x}}_{0}),\cdots,F({\bm{x}}_{t-1})).$
Therefore, we can apply Theorem E.1 to complete the proof. ∎
Appendix F The F2SA-2 Algorithm
We present the realization of F2SA-$p$ when $p=2$ in Algorithm 3 to further compare its procedure with the original F2SA algorithm. Let $\lambda=1/\nu$. We can observe that F2SA [28, 10] solves the following asymmetric penalty problem
$\min_{{\bm{x}}\in{\mathbb{R}}^{d_{x}},{\bm{y}}\in{\mathbb{R}}^{d_{y}}}f({\bm{x}},{\bm{y}})+\lambda\left(g({\bm{x}},{\bm{y}})-\min_{{\bm{z}}\in{\mathbb{R}}^{d_{y}}}g({\bm{x}},{\bm{z}})\right),$
while F2SA-2 solved the following symmetric penalty problem:
$\min_{{\bm{x}}\in{\mathbb{R}}^{d_{x}},{\bm{y}}\in{\mathbb{R}}^{d_{y}}}\frac{1}{2}\left(f({\bm{x}},{\bm{y}})+\lambda f({\bm{x}},{\bm{y}})-\min_{{\bm{z}}\in{\mathbb{R}}^{d_{y}}}(-f({\bm{x}},{\bm{z}})+\lambda g({\bm{x}},{\bm{z}})))\right).$
The latter is better since the symmetric form makes the first-order approximation error to $\nabla\varphi({\bm{x}})$ perfectly cancel out and leave only the second-order error term. Therefore, in terms of the theoretical guarantee by Theorem 3.1, the $\tilde{\mathcal{O}}(\epsilon^{-5})$ upper bound of F2SA-2 can improve the $\tilde{\mathcal{O}}(\epsilon^{-6})$ upper bound of F2SA by a factor of $\epsilon^{-1}$.
Algorithm 3 F2SA-2 $({\bm{x}}_{0},{\bm{y}}_{0})$
[…]
2:for $t=0,1,\cdots,T-1$
3:  Sample random i.i.d indexes $(\xi^{y},\zeta^{y})$, $(\xi^{z},\zeta^{z})$, and $\{({\xi_{i}^{x}},{\zeta_{i}^{x}})\}_{i=1}^{S}.$
4:  ${\bm{y}}_{t}^{0}={\bm{y}}_{t},~{\bm{z}}_{t}^{0}={\bm{z}}_{t}$
5:  for $k=0,1,\cdots,K-1$
6:   ${\bm{y}}_{t}^{k+1}={\bm{y}}_{t}^{k}-\eta_{y}\left(\nu F_{y}({\bm{x}}_{t},{\bm{y}}_{t}^{k};\xi^{y})+G_{y}({\bm{x}}_{t},{\bm{y}}_{t}^{k};\zeta^{y})\right)$
7:   ${\bm{z}}_{t}^{k+1}={\bm{z}}_{t}^{k}-\eta_{y}\left(-\nu F_{y}({\bm{x}}_{t},{\bm{z}}_{t}^{t};\xi^{z})+G_{y}({\bm{x}}_{t},{\bm{z}}_{t}^{k};\zeta^{z})\right)$
8:  end for
9:  ${\bm{y}}_{t+1}={\bm{y}}_{t}^{K},~{\bm{z}}_{t+1}={\bm{z}}_{t}^{K}$
10:  $\Phi_{t}=\dfrac{1}{2}\sum_{i=1}^{S}\left(F_{x}({\bm{x}}_{t},{\bm{y}}_{t+1};{\xi_{i}^{x}})+F_{x}({\bm{x}}_{t},{\bm{z}}_{t+1};\xi_{i}^{x})+\dfrac{G_{x}({\bm{x}}_{t},{\bm{y}}_{t+1};\zeta_{i}^{x})-G_{x}({\bm{x}}_{t},{\bm{z}}_{t+1};\zeta_{i}^{x})}{\nu}\right)$
11:  ${\bm{x}}_{t+1}={\bm{x}}_{t}-\eta_{x}{\Phi_{t}}/{\|\Phi_{t}\|}$
12:end for